# FalconFS: Distributed File System for Large-Scale Deep Learning Pipeline

### v3

@@ 1 Introduction
Distributed file systems (DFSs) are essential components of modern data centers. By providing POSIX-compliant file interfaces within a unified, hierarchical directory structure, DFSs enable general access to underlying storage resources, thereby simplifying storage management and facilitating data sharing among diverse applications [3, 36]. As a result, DFSs form the foundational storage layer for critical data center services, such as block storage and object storage [28, 25], as well as for a broad array of applications, including data processing and analysis [43, 39], high-performance computing [26, 53], and artificial intelligence pipelines [47, 52, 35].
@@
To mitigate path resolution overhead, existing DFSs employ client-side metadata caching, where clients maintain a local cache of directory/file metadata to avoid frequent remote lookups [38, 49, 4, 14, 47, 37, 26, 28, 35, 15]. We refer to clients with client-side caching as stateful clients, as they maintain metadata state locally. By caching resolved paths, stateful clients reduce network round-trips for previously resolved and locally cached file operations, improving performance.
@@ 2.1 Deep Learning Pipeline
Fig. 3 illustrates the architecture of the deep learning (DL) training pipeline employed in Huawei’s autonomous driving (AD) systems. The pipeline consists of four key stages: The ingestion stage collects raw data from real-world environments. The labeling stage generates labels for the raw data with a sequence of model inference tasks, including moving objects detection, lane detection, traffic sign detection, etc. Then, the training stage uses a labeled dataset to train the target model for vehicle deployment. In the final archiving stage, the dataset is moved into low-cost storage systems such as cloud data lakes for future reference [50, 31, 52, 16]. Similar architectures also exist in other DL workloads [52, 2, 32].
@@ 2.2 Workload Patterns in DL pipelines
Enumerous small objects in large directories.
@@
CPU and memory are scarce resources for computing nodes. Training tasks perform data augmentation on CPUs, consuming significant CPU cycles and memory resources. In production, CPU for data augmentation is often the bottleneck, and it is typical to store and reuse intermediate results in memory to reduce CPU load [16, 22]. Therefore, the resources available to DFS clients are limited.
@@ 2.3 Challenge 1: Lookup Tax
For small-file intensive workloads like DL pipelines, metadata performance often becomes the bottleneck. To accelerate metadata operations, most DFSs employ a stateful client architechture that caches metadata on the client side [38, 49, 4, 14, 47, 37, 26, 28, 35, 15]. However, this approach is inefficient for DL training workloads, as their large directory working sets are inefficient to cache, leading to the following dilemma.
@@ 2.4 Challenge 2: Transient Skewness
We observe that when the directory size exceeds the IO parallelism of the tasks, the DFS’s performance degrades. Analysis of load distribution (Fig. 4) shows that at a directory size of 100, the MDSs experience severe load imbalance during read operations. This issue arises because CephFS tends to store metadata for files within the same directory together; consequently, burst operations on files in the same directory can congest a single MDS, leading to performance bottlenecks. Similar patterns also exist in other DFSs like [15, 37, 28, 47, 35].
@@ 3 Proposal: DFS with Stateless Client
Given the challenges brought by DL workloads, we propose the stateless-client architecture for DFS under DL workloads. The stateless-client architecture abandons the cache on the client side and moves path resolution to the server side. The proposal is based on the following three observations. (1) A typical DL cluster usually has far more clients than metadata servers. In Huawei’s deployment environment, the ratio of clients to servers exceeds 40:1. Considering that all clients share the same dataset in DL pipelines, a server-side dentry can serve more clients than a client-side cache. (2) Stateful-client DFSs typically use VFS dcache and inode cache to cache directory attributes [49, 4, 1, 15], which takes 800 bytes for each directory. On the server side, a directory entry can be maintained in a custom format that takes less than 100 bytes. (3) The server has more memory resources than the clients due to the tight resource budget of the computing node, as we described in §2.2. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain.
@@
  • Path-walk indexing. Most DFSs use indexing methods that are related to the parent directories, for example, using the parent directory ID as (part of) the key [26, 15, 28, 47, 35] for hashing, or indexing files in different directories with different hashing rules [14, 37, 4]. These approaches require a full-path walk to resolve inode location, and need caching directory entries on clients to achieve better performance, suffering the trade-off in §2.3.
  • Full-path hashing [40, 42] determines the target inode location by hashing the full path. It does not require cache, but makes directory rename prohibitively expensive due to the relocation of all inodes in the subtree.
@@ 4.1 FalconFS Overview
Metadata nodes (MNode) are PostgreSQL databases [7] with customized extensions. We manage metadata for FalconFS in the extensions using the database’s table and transaction management, B-link tree index, xlog (write-ahead logging), and primary-secondary replication. Each MNode maintains a lazily synchronized directory tree structure (namespace replica) and a partition of file attributes (inodes). Leveraging the merits of central location as servers, MNode merges concurrent requests (§4.4) to de-duplicate shareable execution processes for higher throughput.
@@ 4.2.1 Hybrid Indexing Methods
Fortunately, we find that DL workloads’ large directory size facilitates the filename hashing, significantly reducing the occurrence of load imbalance. Specifically, DL datasets usually have a large directory size, which can be multiple times larger than the number of MNodes. We analyze Huawei’s in-production datasets and popular open-sourced datasets that contain more than ten thousand files [17, 13, 11, 27, 45], finding their directory size ranging from several hundred to hundreds of thousands. According to the law of large numbers, given such large directory sizes, the distribution of files in each directory to MNodes is likely to be uniform. Consequently, the entire namespace is statistically uniformly distributed as the superposition of the distributions for each directory. This forms the basis of using filename hashing as the indexing method for the common cases in DL workloads.
@@ 4.3 Lazy Namespace Replication
Lazy namespace replication is another key enabler of the stateless-client architecture. FalconFS maintains a consistent but not necessarily complete namespace replica on each MNode and the coordinator, enabling local path resolution. To reduce the overhead of maintaining consistency across all replicas, modifications to the namespace are lazily synchronized, and an invalidation-based mechanism is adopted, inspired by cache coherence protocols and Hermes [20]. The design is guided by two principles:
@@ 4.4 Concurrent Request Merging
During path resolution, the worker thread acquires shared locks for all directories along the path to maintain path validity during operations, akin to the implementation of VFS and existing DFS. Prior research demonstrates that lock overhead can be significant even without actual blocking [47]. FalconFS mitigates this through lock coalescing, combining lock acquisition and release operations at per-batch granularity to reduce overhead.
@@
Operations such as mkdir, create, and close modify the inode table. To maintain file system metadata consistency, DFSs typically warp each operation into a separate transaction to persist in atomic [34, 47, 28]. When a transaction commits, it synchronously appends the write-ahead log, leading to small writes that are unfriendly for storage. In FalconFS, as concurrent operations are batched into a single transaction, the worker coalesces small log appends into larger ones, improving the storage efficiency.
@@ 5 Implementation of VFS Compatibility
We observe that the existing semantics provided to the lookup() method are sufficient to distinguish lookup intentions. Since Linux kernel 5.7, the VFS sets the global state flag LOOKUP_PARENT during path walk to indicate that the final component has not yet been reached — a feature designed initially for the kernel audit subsystem [6], and the flag is passed to the lookup() method. If the flag is set, the client module knows that the lookup is for an intermediate directory and returns fake attributes (e.g., $mode=0777$, along with special uiduid and gidgid values) to pass VFS checking.
@@ 6.1 Environment Setup
We compare FalconFS with CephFS 12.2.13 [49], JuiceFS 1.2.1 [1] and Lustre 2.15.6 [4]. CephFS is a widely deployed DFS in data centers. JuiceFS is an open-source DFS targeting AI and data analytics workloads, and we deploy it with TiKV 1.16.1 [8] as its metadata engine and data storage. Lustre is a high-performance DFS widely used in HPC and data centers. CephFS is accessed via the libcephfs library due to observed instability in performance when using a VFS mount point. JuiceFS and Lustre are accessed through VFS mount points, as is FalconFS unless stated otherwise. FalconFS utilizes a modified FUSE kernel module and library that incorporates the optimizations described in §5. All DFSs disable metadata and data replication.
@@ 6.6 Load Balance in Real Workloads
ImageNet [17] 2027728 6.29% 6.21% 0 0
KITTI [13] 15003 7.01% 5.47% 0 0
Cityscapes [11] 20022 6.30% 6.22% 0 0
CelebA [27] 202599 6.54% 6.95% 0 0
SVHN [33] 33404 6.77% 5.76% 0 0
CUB-200-2011 [45] 12003 6.68% 5.95% 0 0
@@
FSL homes [41] 655177 6.83% 5.45% 1 0
In this section, we demonstrate that hybrid metadata indexing achieves a balanced distribution of inodes across diverse directory structures, with only a small portion of filenames requiring special treatment. We evaluate inode distribution on both DL workloads and general-purpose workloads. For DL workloads, we analyze a dataset collected from Huawei’s production environment, as well as six popular open-source image datasets used in deep learning. We select these open-source datasets by listing the most popular image datasets on a dataset summary website [5], and choose the first six datasets containing more than 10,000 files. For general-purpose workloads, we select the Linux 6.8 source tree and FSL home traces [41], the latter being a snapshot of students’ home directories from a shared NFS on a university campus.
@@ 7 Related Works
Path resolution overhead has drawn research attention for a long time. In the context of local file systems, a series of studies propose optimizations like full-path indexing [23, 44, 10] and VFS modifications [48]. These approaches optimize local data structures and cannot be directly applied to distributed file systems (DFSs).
To accelerate path resolution, DFSs typically adopt client-side metadata caching [38, 37, 26, 49, 4, 15, 28]. InfiniFS [28] reduces the cache misses penalty by resolving multiple path components in parallel; however, it cannot mitigate request amplification. Giraffa [40] and CalvinFS [42] locate inodes by full path hashing, which makes directory renaming hard to implement. HDFS [39] performs path resolution on a centralized namenode and thus has scalability issues. HopsFS [34] performs path resolution at a proxy layer, which looks up all path components in parallel from a distributed database, but still suffers from constant request amplification. Our approach differs in that FalconFS addresses scalability, request amplification, and metadata indexing issues through a client-stateless architecture and hybrid metadata indexing. Mantle [24], a concurrent work of FalconFS, adopts a similar client-stateless metadata service to object storage services. While FalconFS scales out path resolution to all NModes, Mantle resolves path in a per-namespace single node and has scalability issues.
@@
Previous studies propose various data loading frameworks for DL training tasks [19, 22, 30, 21, 51]. These works optimize DL data loading by unifying data access, reusing pre-processed data, and leveraging data caching, etc. These works are task-specific and can be deployed on top of FalconFS. DIESEL [46] is a DFS designed for DL training tasks. Its design is based on the assumption that the dataset’s directory tree is small enough to be cached on every client, whereas FalconFS makes the opposite assumption and focuses on eliminating client-side caching. 3FS [9] is a recent DFS for AI workloads. Unlike FalconFS, it is optimized for large data access and focuses on optimizing the data path, while FalconFS optimizes the metadata architecture for small-file performance.
@@ Acknowledgements
We sincerely thank our shepherd, Ken Birman, and the anonymous reviewers for their constructive comments and insightful suggestions. This work is supported in part by the National Natural Science Foundation of China (No. 62132014), the Fundamental Research Funds for the Central Universities, and Huawei Technologies. Mingkai Dong (mingkaidong@sjtu.edu.cn) and Junbin Kang (kangjunbin1@huawei.com) are the corresponding authors.
@@ A.1 Theoretical Analysis
Our proof builds upon a theorem from caching literature [12, 18], which states: for $m$ objects randomly partitioned across $n$ nodes with a total query load of $n\cdot t$, if a cache absorbs all queries to the hottest $O(nlogn)$ items, then no node exceeds $t$ load with high probability, independent of the query distribution. We adapt this through constructive proof.

### v4

@@ 1 Introduction
Distributed file systems (DFSs) are essential components of modern data centers. By providing POSIX-compliant file interfaces within a unified, hierarchical directory structure, DFSs enable general access to underlying storage resources, thereby simplifying storage management and facilitating data sharing among diverse applications [3FS, GeminiFS]. As a result, DFSs form the foundational storage layer for critical data center services, such as block storage and object storage [Lv2022InfiniFS, 285788], as well as for a broad array of applications, including data processing and analysis [Hive, Shvachko2010The-Hadoop], high-performance computing [Li2017LocoFS, DeltaFS], and artificial intelligence pipelines [CFS, Zhao2022Understanding, Pan2021Facebooktextquoterights].
@@
To mitigate path resolution overhead, existing DFSs employ client-side metadata caching, where clients maintain a local cache of directory/file metadata to avoid frequent remote lookups [nfsv4, Weil2006Ceph, lustre, 2019Gluster, CFS, Ren2014IndexFS, Li2017LocoFS, Lv2022InfiniFS, Pan2021Facebooktextquoterights, BeeGFS]. We refer to clients with client-side caching as stateful clients, as they maintain metadata state locally. By caching resolved paths, stateful clients reduce network round-trips for previously resolved and locally cached file operations, improving performance.
@@ 2.1 Deep Learning Pipeline
Fig. 3 illustrates the architecture of the deep learning (DL) training pipeline employed in Huawei’s autonomous driving (AD) systems. The pipeline consists of four key stages: The ingestion stage collects raw data from real-world environments. The labeling stage generates labels for the raw data with a sequence of model inference tasks, including moving objects detection, lane detection, traffic sign detection, etc. Then, the training stage uses a labeled dataset to train the target model for vehicle deployment. In the final archiving stage, the dataset is moved into low-cost storage systems such as cloud data lakes for future reference [Zaharia2021LakehouseAN, tf.data, Zhao2022Understanding, Cachew]. Similar architectures also exist in other DL workloads [Zhao2022Understanding, netapp-dp, netapp-ad].
@@ 2.2 Workload Patterns in DL pipelines
Numerous small objects in large directories.
@@
CPU and memory are scarce resources for computing nodes. Training tasks perform data augmentation on CPUs, consuming significant CPU cycles and memory resources. In production, CPU for data augmentation is often the bottleneck, and it is typical to store and reuse intermediate results in memory to reduce CPU load [Cachew, Revamper]. Therefore, the resources available to DFS clients are limited.
@@ 2.3 Challenge 1: Lookup Tax
For small-file intensive workloads like DL pipelines, metadata performance often becomes the bottleneck. To accelerate metadata operations, most DFSs employ a stateful client architechture that caches metadata on the client side [nfsv4, Weil2006Ceph, lustre, 2019Gluster, CFS, Ren2014IndexFS, Li2017LocoFS, Lv2022InfiniFS, Pan2021Facebooktextquoterights, BeeGFS]. However, this approach is inefficient for DL training workloads, as their large directory working sets are inefficient to cache, leading to the following dilemma.
@@ 2.4 Challenge 2: Transient Skewness
We observe that when the directory size exceeds the IO parallelism of the tasks, the DFS’s performance degrades. Analysis of load distribution (Fig. 4) shows that at a directory size of 100, the MDSs experience severe load imbalance during read operations. This issue arises because CephFS tends to store metadata for files within the same directory together; consequently, burst operations on files in the same directory can congest a single MDS, leading to performance bottlenecks. Similar patterns also exist in other DFSs like [BeeGFS, Ren2014IndexFS, Lv2022InfiniFS, CFS, Pan2021Facebooktextquoterights].
@@ 3 Proposal: DFS with Stateless Client
Given the challenges brought by DL workloads, we propose the stateless-client architecture for DFS under DL workloads. The stateless-client architecture abandons the cache on the client side and moves path resolution to the server side. The proposal is based on the following three observations. (1) A typical DL cluster usually has far more clients than metadata servers. In Huawei’s deployment environment, the ratio of clients to servers exceeds 40:1. Considering that all clients share the same dataset in DL pipelines, a server-side dentry can serve more clients than a client-side cache. (2) Stateful-client DFSs typically use VFS dcache and inode cache to cache directory attributes [Weil2006Ceph, lustre, juicefs, BeeGFS], which takes 800 bytes for each directory. On the server side, a directory entry can be maintained in a custom format that takes less than 100 bytes. (3) The server has more memory resources than the clients due to the tight resource budget of the computing node, as we described in §2.2. However, to realize the stateless-client architecture for an efficient DFS, two key problems remain.
@@
  • Path-walk indexing. Most DFSs use indexing methods that are related to the parent directories, for example, using the parent directory ID as (part of) the key [Li2017LocoFS, BeeGFS, Lv2022InfiniFS, CFS, Pan2021Facebooktextquoterights] for hashing, or indexing files in different directories with different hashing rules [2019Gluster, Ren2014IndexFS, lustre]. These approaches require a full-path walk to resolve inode location, and need caching directory entries on clients to achieve better performance, suffering the trade-off in §2.3.
  • Full-path hashing [Giraffa, CalvinFS] determines the target inode location by hashing the full path. It does not require cache, but makes directory rename prohibitively expensive due to the relocation of all inodes in the subtree.
@@ 4.1 FalconFS Overview
Metadata nodes (MNode) are PostgreSQL databases [PostgreSQL] with customized extensions. We manage metadata for FalconFS in the extensions using the database’s table and transaction management, B-link tree index, xlog (write-ahead logging), and primary-secondary replication. Each MNode maintains a lazily synchronized directory tree structure (namespace replica) and a partition of file attributes (inodes). Leveraging the merits of central location as servers, MNode merges concurrent requests (§4.4) to de-duplicate shareable execution processes for higher throughput.
@@ 4.2.1 Hybrid Indexing Methods
Fortunately, we find that DL workloads’ large directory size facilitates the filename hashing, significantly reducing the occurrence of load imbalance. Specifically, DL datasets usually have a large directory size, which can be multiple times larger than the number of MNodes. We analyze Huawei’s in-production datasets and popular open-sourced datasets that contain more than ten thousand files [imagenet-object-localization-challenge, Geiger2012CVPR, Cordts2016Cityscapes, liu2015faceattributes, WahCUB_200_2011], finding their directory size ranging from several hundred to hundreds of thousands. According to the law of large numbers, given such large directory sizes, the distribution of files in each directory to MNodes is likely to be uniform. Consequently, the entire namespace is statistically uniformly distributed as the superposition of the distributions for each directory. This forms the basis of using filename hashing as the indexing method for the common cases in DL workloads.
@@ 4.3 Lazy Namespace Replication
Lazy namespace replication is another key enabler of the stateless-client architecture. FalconFS maintains a consistent but not necessarily complete namespace replica on each MNode and the coordinator, enabling local path resolution. To reduce the overhead of maintaining consistency across all replicas, modifications to the namespace are lazily synchronized, and an invalidation-based mechanism is adopted, inspired by cache coherence protocols and Hermes [Hermes]. The design is guided by two principles:
@@ 4.4 Concurrent Request Merging
During path resolution, the worker thread acquires shared locks for all directories along the path to maintain path validity during operations, akin to the implementation of VFS and existing DFS. Prior research demonstrates that lock overhead can be significant even without actual blocking [CFS]. FalconFS mitigates this through lock coalescing, combining lock acquisition and release operations at per-batch granularity to reduce overhead.
@@
Operations such as mkdir, create, and close modify the inode table. To maintain file system metadata consistency, DFSs typically warp each operation into a separate transaction to persist in atomic [Niazi2017HopsFS, CFS, Lv2022InfiniFS]. When a transaction commits, it synchronously appends the write-ahead log, leading to small writes that are unfriendly for storage. In FalconFS, as concurrent operations are batched into a single transaction, the worker coalesces small log appends into larger ones, improving the storage efficiency.
@@ 5 Implementation of VFS Compatibility
We observe that the existing semantics provided to the lookup() method are sufficient to distinguish lookup intentions. Since Linux kernel 5.7, the VFS sets the global state flag LOOKUP_PARENT during path walk to indicate that the final component has not yet been reached — a feature designed initially for the kernel audit subsystem [path-lookup], and the flag is passed to the lookup() method. If the flag is set, the client module knows that the lookup is for an intermediate directory and returns fake attributes (e.g., $mode=0777$, along with special uiduid and gidgid values) to pass VFS checking.
@@ 6.1 Environment Setup
We compare FalconFS with CephFS 12.2.13 [Weil2006Ceph], JuiceFS 1.2.1 [juicefs] and Lustre 2.15.6 [lustre]. CephFS is a widely deployed DFS in data centers. JuiceFS is an open-source DFS targeting AI and data analytics workloads, and we deploy it with TiKV 1.16.1 [TiKV] as its metadata engine and data storage. Lustre is a high-performance DFS widely used in HPC and data centers. CephFS is accessed via the libcephfs library due to observed instability in performance when using a VFS mount point. JuiceFS and Lustre are accessed through VFS mount points, as is FalconFS unless stated otherwise. FalconFS utilizes a modified FUSE kernel module and library that incorporates the optimizations described in §5. All DFSs disable metadata and data replication.
@@ 6.6 Load Balance in Real Workloads
ImageNet [imagenet-object-localization-challenge] 2027728 6.29% 6.21% 0 0
KITTI [Geiger2012CVPR] 15003 7.01% 5.47% 0 0
Cityscapes [Cordts2016Cityscapes] 20022 6.30% 6.22% 0 0
CelebA [liu2015faceattributes] 202599 6.54% 6.95% 0 0
SVHN [Netzer2011ReadingDI] 33404 6.77% 5.76% 0 0
CUB-200-2011 [WahCUB_200_2011] 12003 6.68% 5.95% 0 0
@@
FSL homes [snia-trace-static-5228] 655177 6.83% 5.45% 1 0
In this section, we demonstrate that hybrid metadata indexing achieves a balanced distribution of inodes across diverse directory structures, with only a small portion of filenames requiring special treatment. We evaluate inode distribution on both DL workloads and general-purpose workloads. For DL workloads, we analyze a dataset collected from Huawei’s production environment, as well as six popular open-source image datasets used in deep learning. We select these open-source datasets by listing the most popular image datasets on a dataset summary website [paperswithcode], and choose the first six datasets containing more than 10,000 files. For general-purpose workloads, we select the Linux 6.8 source tree and FSL home traces [snia-trace-static-5228], the latter being a snapshot of students’ home directories from a shared NFS on a university campus.
@@ 7 Related Works
Path resolution overhead has drawn research attention for a long time. In the context of local file systems, a series of studies propose optimizations like full-path indexing [10.1145/2485732.2485741, 10.1145/2815400.2815405, FlatFS] and VFS modifications [216892]. These approaches optimize local data structures and cannot be directly applied to distributed file systems (DFSs).
To accelerate path resolution, DFSs typically adopt client-side metadata caching [nfsv4, Ren2014IndexFS, Li2017LocoFS, Weil2006Ceph, lustre, BeeGFS, Lv2022InfiniFS]. InfiniFS [Lv2022InfiniFS] reduces the cache misses penalty by resolving multiple path components in parallel; however, it cannot mitigate request amplification. Giraffa [Giraffa] and CalvinFS [CalvinFS] locate inodes by full path hashing, which makes directory renaming hard to implement. HDFS [Shvachko2010The-Hadoop] performs path resolution on a centralized namenode and thus has scalability issues. HopsFS [Niazi2017HopsFS] performs path resolution at a proxy layer, which looks up all path components in parallel from a distributed database, but still suffers from constant request amplification. Our approach differs in that FalconFS addresses scalability, request amplification, and metadata indexing issues through a client-stateless architecture and hybrid metadata indexing. Mantle [Mantle], a concurrent work of FalconFS, adopts a similar client-stateless metadata service to object storage services. While FalconFS scales out path resolution to all NModes, Mantle resolves path in a per-namespace single node and has scalability issues.
@@
Previous studies propose various data loading frameworks for DL training tasks [OneAccess, Revamper, CoorDL, Quiver, SiloD]. These works optimize DL data loading by unifying data access, reusing pre-processed data, and leveraging data caching, etc. These works are task-specific and can be deployed on top of FalconFS. DIESEL [DIESEL] is a DFS designed for DL training tasks. Its design is based on the assumption that the dataset’s directory tree is small enough to be cached on every client, whereas FalconFS makes the opposite assumption and focuses on eliminating client-side caching. 3FS [Fire-Flyer] is a recent DFS for AI workloads. Unlike FalconFS, it is optimized for large data access and focuses on optimizing the data path, while FalconFS optimizes the metadata architecture for small-file performance.
@@ Acknowledgements
We sincerely thank our shepherd, Ken Birman, and the anonymous reviewers for their constructive comments and insightful suggestions. This work is supported in part by the National Natural Science Foundation of China (No. 62132014), the Fundamental Research Funds for the Central Universities, the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (JYB2025XDXM113), and Huawei Technologies. Mingkai Dong (mingkaidong@sjtu.edu.cn) and Junbin Kang (kangjunbin1@huawei.com) are the corresponding authors.
@@ A.1 Theoretical Analysis
Our proof builds upon a theorem from caching literature [10.1145/2038916.2038939, Jin2017NetCache], which states: for $m$ objects randomly partitioned across $n$ nodes with a total query load of $n\cdot t$, if a cache absorbs all queries to the hottest $O(nlogn)$ items, then no node exceeds $t$ load with high probability, independent of the query distribution. We adapt this through constructive proof.